# Granger Causality in High-Dimensional Networks of Time Series

### v2

@@ Abstract
A novel approach is developed for discovering directed connectivity between specified pairs of nodes in a high-dimensional network (HDN) of brain signals. To accurately identify causal connectivity for such specified objectives, it is necessary to properly address the influence of all other nodes within the network. The proposed procedure herein starts with the estimation of a low-dimensional representation of the other nodes in the network utilizing (frequency-domain-based) spectral dynamic principal component analysis (sDPCA). The resulting scores can then be removed from the nodes of interest, thus eliminating the confounding effect of other nodes within the HDN. Accordingly, causal interactions can be dissected between nodes that are isolated from the effects of the network. Extensive simulations have demonstrated the effectiveness of this approach as a tool for causality analysis in complex time series networks. The proposed methodology has also been shown to be applicable to multichannel EEG networks.
Index Terms:
@@
High-Dimensional Networks of Signals, Connectivity, Causality, Spectral Dynamic PCA, EEG signals
I Introduction
Identifying effective connectivity in high-dimensional networks (HDN) is a challenging problem that remains an active area of research and has been previously addressed by several studies [see, e.g., 1, 2, 3, 4, 5, 6, 7, 8]. Moreover, it is difficult to infer or interpret interactions between a pair of nodes while taking into account potential interactions with all other nodes in the background [9, 10]. This is a particularly relevant issue (e.g., analysis of EEG-like signals) when the focus might stand on a few channels/nodes of interest (COI/NOI) within an extensive network of interacting nodes. In fact, utilizing conventional vector autoregressive models (VAR) can be a practical way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed. In this paper, we propose a framework that provides a low-dimensional summary of the background activity in the high-dimensional network (HDN).
To discern effective connectivity, particularly between pairs of nodes in an HDN of brain signals, we propose a practical and easily accessible approach based on spectral dynamic principal component analysis (PCA) ([11, 12]) combined with testing for GC in lieu of the conventional F-test.
@@ I Introduction
The sDPCA method enhances the classical PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining maximal temporal information. In other words, while the classical PCA is based on the eigendecomposition of the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral matrix at all frequencies. Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags) then this approach captures lead-lag relationships amongst signals.
Various methodologies have evolved from sDPCA. For instance, [13] developed a spectral envelope using sDPCA to identify common signals across multiple time series, while [14] proposed a functional variant of sDPCA for analyzing functional time-series data. As indicated by [15, 5, 16], spectral principal component analysis (sPCA) in the frequency domain offers practical benefits and is highly effective in low-dimensional representations of high-dimensional time series. The main objective of this study is to illustrate how the proposed approach enables the practical and sufficient isolation of COI by partialling out the influence of disinterested nodes, which are derived as a dynamic summary in a low-dimensional representation. The frequency-domain dynamic PCA method maintains the spectral characteristics of the data, which are crucial for accurately determining causal relationships in time series networks.
The proposed approach offers two main advantages. Firstly, sDPCA in the frequency domain offers an accurate and practical dimension reduction while preserving the spectral information. This is essential because the spectrum of a signal captures cross-covariances at all lags. This leads to a more precise isolation of COI, reducing the likelihood of spurious causal inferences. Secondly, the proposed approach enables the formation of a comprehensive yet interpretable representation of the interactions between specified nodes at a defined level of resolution rather than being unable to infer by the expansive number of possible interactions between nodes. This facilitates the study of considerable numbers of interactions, ranging from pairs of nodes to interactions between a tractable number of clusters of nodes. The subsequent sections will elaborate on the proposed approach and its components. Section 2 presents the proposed approach. Section 3 outlines the scope of the simulation study and presents related results. Section 4 applies the proposed approach to EEG data and discusses the findings.
II Methodology
II-A Inference in high-dimensional networks of time series
@@
Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of signals with possibly $n\geq 256$ nodes. As we consider in this paper, this network may represent the activity of a brain through EEG recordings in which each component (i.e., node or channel) is actually a time series. In this notation of network $N$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest. For problems whose solutions are sought component-wise, we set our goal as exploring the interaction or efficient connectivity between any pair of nodes in a network rather than the relationality of the whole network. These two nodes are subject to the influence of other nodes due to the network’s relational structure. This means that to dissect the respective efficient connection between any two nodes of interest; they need to be freed from the influences of other nodes. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
$Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
$X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
@@
where the $f_{z}(.)$ is a function that encompasses the dynamic variability of excluded nodes. Practically, the function $f_{z}(.)$ acts as a filter and can take any form as long as it provides the net summary discharge of the nodes excluded. Within the scope of this study, we utilize the spectral domain dynamic PCA implementation made available by Hörmann et al. (2015) via R package called freqdom. This method, originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence [17, 16, see, e.g., ]. Hence, we propose the filter $f_{z}(.)$ over excluded nodes to provide the following summary indicators,
$f_{z}(Z_{1,t},\ldots,Z_{n-2,t})=\{pc1,pc2,\ldots,pc_{k},\text{interactions}\},$
where $pc1,pc2,\ldots,pc_{k}$ are the scores of the principal components obtained through sDPCA, and interactions represent any interaction terms between these principal components.
The conditional means with respect to these indicators are then given by:
$E\big{(}Y^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=a_{1}\cdot pc1+a_{2}\cdot pc+$
$\cdots+a_{k}\cdot pc_{k}+\sum a_{ij}\cdot(pc_{i}\times pc_{j}),$
$E\big{(}X^{*}_{t}\mid f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)}=b_{1}\cdot pc1+b_{2}\cdot pc2+$
$\cdots+b_{k}\cdot pc_{k}+\sum b_{ij}\cdot(pc_{i}\times pc_{j}),$
where $pc_{i}$ represents the score of the $i$-th dynamic principal component and $a_{i},b_{i},a_{ij},b_{ij}$ are coefficients corresponding to the principal components and their interactions. The number of principal components $k$ can be chosen to be relatively low compared to the number of dimensions, provided that they capture a sufficient proportion of the total variability in the network. For practical implementation, the first few principal components are generally sufficient.
@@
The dynamic principal component scores $dpc_{t}^{(j)}$ are then calculated by filtering the time series $\mathbf{Z}_{t}$ with the dynamic principal component filters:
$dpc_{t}^{(j)}=\sum_{k=-q}^{q}\phi_{k}^{(j)}\mathbf{Z}_{t-k}.$
This convolution operation can be represented in matrix notation as:
$\mathbf{dpc}=\mathbf{Z}\ast\mathbf{\Phi},$
where $\ast$ denotes the convolution operation.
@@
Terms and Notation
  • $\mathbf{Z}$: Original multivariate time series matrix of dimensions $T\times n$, where $T$ is the number of time points and $n$ is the number of variables.
  • $F_{\mathbf{Z}}(\omega)$: Cross-spectral density matrix at frequency $\omega$.
  • $\text{Cov}(\mathbf{Z}_{h},\mathbf{Z}_{0})$: Covariance matrix at lag $h$.
  • $\hat{C}_{\mathbf{Z}}(h)$: Empirical lagged auto-covariance matrix at lag $h$.
  • $w(\cdot)$: Window function used to smooth the estimates.
  • $q$: Window size for the kernel estimator.
  • $\phi_{k}^{(j)}$: Dynamic principal component filters, obtained as the Fourier coefficients of the dynamic eigenvectors.
  • $\varphi_{j}(\omega)$: Dynamic eigenvectors of the spectral density matrix $F(\omega)$.
  • $dpc_{t}^{(j)}$: Dynamic principal component scores.
  • $\mathbf{dpc}$: Matrix of dynamic principal component scores.
  • $\mathbf{\Phi}$: Matrix of dynamic principal component filters.
  • $\lambda_{j}(\omega)$: $j$-th dynamic eigenvalue of the spectral density matrix $F(\omega)$.
II-B Granger Causality test
Granger Causality (GC) is a statistical hypothesis test used to determine whether one time series can predict another. The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to ”Granger-cause” $Y_{t}$.
The Granger causality test involves the following steps. First, consider two time series $X_{t}$ and $Y_{t}$. To test if $X_{t}$ Granger-causes $Y_{t}$, we compare the following two models:
@@ II-B Granger Causality test
II-C Proposed Approach
In high-dimensional networks of time series, analyzing causal connectivity between specific pairs of signals can be both challenging and computationally intensive due to the complexity of modeling interactions across all nodes. The approach to be followed would be to model the whole material if there were no high dimensionality problems. However, modeling under high dimensionality requires fitting highly parameterized models to the entire network, leading to difficult-to-interpret estimates and an increased risk of overfitting. To overcome these challenges, we propose a practical and easily implementable method that focuses on isolating the nodes/channels of interest by effectively reducing the dimensionality of the network. Our approach utilizes spectral dynamic principal component analysis to capture the essential variability of the background network in a low-dimensional representation. By partialling out the influence of other nodes through sDPCA, we can isolate the COI and accurately assess the causal connectivity between them using conventional Granger causality tests. This method simplifies the analysis by avoiding the need to model the entire high-dimensional network and enhances interpretability when the primary interest is in the causal relationship between specific signals. The proposed approach is given in the Procedure outline below.
Input: High-dimensional Networks of Signals $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
@@ II-C Proposed Approach
1. Identify Nodes of Interest: NOI:$\{Y^{*}_{t},X^{*}_{t}\}$.
In practice, each node can also be a cluster of nodes that can be represented by sDPCA scores.
@@ 2. Apply sDPCA:
    Estimate spectral density matrix $F_{\mathbf{Z}}(\omega)$ for $\mathbf{Z}_{t}=\{Z_{1,t},\ldots,Z_{n-2,t}\}$.
@@
    Compute dynamic eigenvectors $\varphi_{j}(\omega)$ and
eigenvalues $\lambda_{j}(\omega)$.
  • iii
    Derive dynamic principal component filters:
[…]
  • iv
    Calculate dynamic principal component scores:
    $\text{dpc}_{t}^{(j)}=\sum_{k=-q}^{q}\phi_{k}^{(j)}\mathbf{Z}_{t-k}$
  • 3. Partial Out Influence of Other Nodes:
    $Y_{t}=Y^{*}_{t}-\hat{E}\big{(}Y^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$
    $X_{t}=X^{*}_{t}-\hat{E}\big{(}X^{*}_{t}\mid\text{dpc}_{t}^{(1)},\ldots,\text{dpc}_{t}^{(k)}\big{)}$

### v3

@@ Abstract
A novel and practical approach is introduced for identifying causal connectivity between specific node pairs in high-dimensional networks (HDN) of brain signals. The novelty lies in enabling a conventional Granger causality analysis to work effectively in a high-dimensional setting by combining it with frequency-domain-based spectral dynamic principal component analysis (sDPCA). The proposed procedure begins with the estimation of a low-dimensional representation of the other nodes in the network utilizing sDPCA. The resulting scores can then be extracted from the nodes of interest, thus removing the confounding effect of other nodes within the HDN. Accordingly, causal interactions can be dissected between nodes that are isolated from the effects of the network. Simulations confirm that the approach reliably retrieves causal links, even in highly interactive networks. The method also performs well on multichannel EEG data, and findings align with earlier findings in the literature.
@@
Keywords: High-Dimensional Networks of Signals, Connectivity, Causality, Spectral Dynamic PCA, EEG signals
1 Introduction
Identifying effective connectivity in high-dimensional networks (HDN) remains a demanding problem and has been addressed by several studies (see, e.g., Kalisch and Bühlmann, 2014; Wang et al., 2016; Ting et al., 2016, 2017; Wang et al., 2019; Zarghami and Friston, 2020; Siggiridou and Kugiumtzis, 2021; Shojaie and Fox, 2022). Moreover, it is difficult to infer or interpret interactions between a pair of nodes while taking into account potential interactions with all other nodes in the background (Wang and Drton, 2020; Basu et al., 2024). One particular challenge arises when the objective is simply to identify the causal interaction between specifically interested pairs of nodes or pairs of subsets of nodes/channels within an extensive network of interacting nodes. Attempting to model an entire HDN for such a specific objective will require fitting a highly parameterized model with potentially hundreds of parameters, which can lead to over-parameterization of the task and interpretation difficulties.
In fact, utilizing conventional bivariate vector autoregressive models (VAR) can be a practical and manageable way to identify pairwise causal interactions in high-dimensional networks as long as the potential confounding effect due to the high-dimensional background nodes is addressed. Adopting this strategy, we propose a procedure that uses conventional Granger causality (GC) testing to detect causal connections between any two nodes while accounting for the global background activity in a low-dimensional representation. Such a representation is constructed through spectral dynamic principal component analysis (sDPCA) (Brillinger, 1969, 1981), which extends the principal component analysis to incorporate temporal dependence. By working with the spectral density matrix —directly linked to all auto and cross-covariances- sDPCA preserves frequency-specific features critical for neural data analysis.
@@ 1 Introduction
The sDPCA enhances the PCA by incorporating temporal dependency within the data, thereby reducing the dimensionality of time series in the frequency domain while retaining temporal information. In other words, while the conventional PCA is based on the eigendecomposition at the zero-lag (or contemporaneous) covariance matrix, sDPCA is based on the eigendecomposition of the spectral density matrix estimated for lagged auto-covariances at all frequencies. Since there is a 1-1 relationship between the spectral matrix (at all frequencies) and the covariance matrix (at all lags), this approach captures lagged cross-correlations amongst signals.
Various methodologies have been derived from frequency domain PCA. For instance, Stoffer et al. (1993); Stoffer (1999); Stoffer et al. (2000) developed a spectral envelope using dynamic PCA to identify common signals in multiple categorical and real-valued time series, Hörmann et al. (2015) proposed a variant of dynamic PCA for analyzing functional time-series data. As demonstrated by Ombao and Ho (2006); Wang et al. (2019); Ombao and Pinto (2022), principal component analysis in the frequency domain offers practical benefits and is highly effective in summarizing high-dimensional signals by capturing their main modes of variation. On the other hand, given the multivariate nature of the objectives outlined herein, it becomes relevant to inquire about the applicability of the state-of-the-art techniques of multivariate time series analysis. For example, Factor-augmented VAR (FAVAR) models (Bernanke et al., 2005; Lin and Michailidis, 2020) reduce dimensionality by extracting a handful of static factors from large panels and then placing these factors into a more manageable VAR system. This strategy centers on global modeling and treating the extracted factors as regressors to explain multiple target variables; however, it is not designated for cases where attention is on a single pair of nodes within a high-dimensional network. Dynamic factor models (DFMs) (Forni et al., 2000, 2015; Stock and Watson, 2005; Barigozzi and Hallin, 2017) similarly rely on a few latent components, focusing on decomposing each series into common and idiosyncratic components. DFM allows more detailed lagged structures than static factor approaches but often seeks to explain broad data variance or perform large-scale forecasts. Conditional Granger causality (CGC) (Geweke, 1984) instead tests whether adding one predictor refines forecasts of another, holding all others fixed. Although CGC reveals causal associations, it can become complicated in high dimensions when many variables must be jointly conditioned. While there are some parallels between the proposed approach and that of others, each method is characterized by a distinct emphasis. The proposed approach diverges by isolating specific node pairs. It accounts for background effects by spectral components and then applies GC testing on a simplified bivariate system. This diverges from CGC’s exhaustive conditioning on many channels, from FAVAR’s global factor plus VAR, and dynamic factor models’ emphasis on large-scale shared patterns. The approach also provides feasibility in large networks, where a smaller factor-driven representation is sufficient to identify causal interactions.
The proposed approach distinguishes itself from existing approaches by focusing on frequency-domain dimensionality reduction combined with targeted pairwise causality analysis. The theoretical formulations of conditional GC, FAVAR, and DFMs have different aims – identifying influences through conditioning, incorporating large information sets into VARs, and explaining common variance in high dimensions, respectively – and each comes with certain assumptions about the data-generating process. The proposed approach does not compete with these state-of-the-art methods on their own terms; rather, it occupies a complementary cavity. It assumes that in many high-dimensional systems (e.g., brain recordings), one can extract the key oscillatory components and concentrate on their interactions, thereby enabling meaningful causal understanding without having to model every other interaction. Empirically, this translates to a convenient method that is sensitive to the features of brain monitoring data – handling many channels and frequency-specific processes – while building on the proven strengths of factor modeling and Granger causality. Given the objectives of established methods and our approach, it is important to note that utilizing GC with sDPCA does not propose a universal substitute for full network modeling. Instead, it is a domain-orientated alternative that emphasizes interpretability and feasibility in the analysis of high-dimensional network time series. This allows for targeted analysis, whether at the level of individual node pairs or among a tractable number of node clusters, without being constrained by finitely many interactions. In what follows, Section 2 presents the proposed approach for an HDN setting, Section 3 describes simulation setups and results, and Section 4 features an application to EEG recordings, showing how the method handles brain signals.
2 Methodology
2.1 Inference in high-dimensional networks of time series
@@
Define $V=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a high dimensional network of signals with possibly $n\geq 256$ nodes. As we consider in this paper, this network may represent the activity of a brain through EEG recordings in which each component (i.e., node or channel) is actually a time series. In this notation of network $V$, $Y^{*}$ and $X^{*}$ denote any two nodes of interest. For problems whose solutions are sought component-wise, we set our goal to explore the effective connectivity between any node pairs in a network rather than the associativity of the whole network. Node pairs are subject to the influence of other nodes due to the network’s relational structure. This means that to dissect the respective effective connection between any two nodes of interest; they need to be freed from the influences of other nodes. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
$\hat{Y}_{t}=Y^{*}_{t}-\hat{Y}^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t}),$
$\hat{X}_{t}=X^{*}_{t}-\hat{X}^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t}),$
@@
where the $f_{z}(.)$ is a function that encompasses the dynamic variability of excluded nodes. Practically, the function $f_{z}(.)$ acts as a filter and can take any form as long as it provides the net summary discharge of the nodes excluded. In the scope of this paper, the function $f_{z}(.)$ is intended to provide low-dimensional summaries of excluded nodes. To this end, we utilize the spectral-domain dynamic PCA implemented in the R package freqdom made available by Hörmann et al. (2015). Spectral PCA originally proposed by Brillinger (1981), enables us to obtain principal component scores that concisely encapsulate variability in multi-dimensional time series, specifically with regard to their serial dependence (Shumway et al., 2000; Ombao and Pinto, 2022, see, e.g., ). Hence, we propose the filter $f_{z}(\cdot)$ over the excluded nodes to yield a vector of summary indicators,
$f_{z}(Z_{1,t},\ldots,Z_{n-2,t})=\mathbf{p}_{t}=\bigl{(}dpc_{1}(t),dpc_{2}(t),\ldots,dpc_{k}(t)\bigr{)},$
where $dpc_{i}(t)$ denotes the score of the $i$-th dynamic principal component obtained through sDPCA. One may optionally augment $\mathbf{p}_{t}$ with higher-order transformations of the principal-component scores (for instance, cross-products $pc_{i}(t)\times pc_{j}(t)$ or polynomial expansions) if a more general functional form is required to capture nonlinearities or interactions among these components.
In general, the conditional mean of $Y^{*}_{t}$ and $X^{*}_{t}$ given these indicators can be modeled as:
$\hat{E}\bigl{(}Y^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}=G_{Y^{*}}\bigl{(}\mathbf{p}_{t};\bm{\theta}_{Y^{*}}\bigr{)},$
$\hat{E}\bigl{(}X^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}=G_{X^{*}}\bigl{(}\mathbf{p}_{t};\bm{\theta}_{X^{*}}\bigr{)},$
where $G_{Y^{*}}(\cdot;\bm{\theta}_{Y^{*}})$ and $G_{X^{*}}(\cdot;\bm{\theta}_{X^{*}})$ are functions (e.g. linear, nonlinear, or another parametric/nonparametric regression form) parameterized by $\bm{\theta}_{Y^{*}}$ and $\bm{\theta}_{X^{*}}$.
Considering an option where the terms are linearly combined and have first-order interactions, the representation can be written as follows:
$G_{Y^{*}}\bigl{(}\mathbf{p}_{t};\bm{\theta}_{Y^{*}}\bigr{)}=\alpha_{0}+\sum_{i=1}^{q}\alpha_{i}dpc_{i}(t)+\sum_{i=1}^{q}\sum_{j\neq i}^{q}\alpha_{ij}\bigl{[}dpc_{i}(t)\times dpc_{j}(t)\bigr{]},$
Here, $q$ is typically small relative to the original dimension, so that $\{\alpha_{0},\alpha_{i},\alpha_{ij}\}$ remain tractable to estimate in practice.
@@
By construction, sDPCA typically employs two-sided filters when forming the scores. This is seen in
$dpc_{t}^{(j)}=\sum_{k=-q}^{q}\bigl{\{}\phi_{k}^{(j)}\bigr{\}}^{\!\top}\mathbf{Z}_{t-k},$
where $\{\phi_{k}^{(j)}\}$ are the Fourier-based filter coefficients arising from the inverse transform of the $j$-th dynamic eigenvector. The index $k$ runs from $-q$ to $q$, reflecting a non-causal (two-sided) convolution that uses both past and future data relative to time $t$. This formulation can yield improved spectral estimates, though it assumes the entire data record is available.
In real-time or causal applications, a one-sided variant can be adopted:
$dpc_{t}^{(j)}=\sum_{k=0}^{q}\bigl{\{}\phi_{k}^{(j)}\bigr{\}}^{\!\top}\mathbf{Z}_{t-k},$
which omits “future” values $\mathbf{Z}_{t+k}$ for $k>0$. This causal filter operates only on current and lagged observations, making it suitable for sequential or forecasting tasks. The choice between two-sided and one-sided implementations depends on whether retrospective or real-time usage is intended, as well as on the degree to which one wishes to exploit possible gains from including both past and future observations.
@@
2.2 Granger Causality
Granger Causality (GC) is a statistical concept used to determine whether one time series can predict another. The concept, introduced by Granger (1969), is based on the principle that if the prediction of a time series $Y_{t}$ can be improved by incorporating past values of another time series $X_{t}$, then $X_{t}$ is said to ”Granger-cause” $Y_{t}$.
The Granger causality analysis involves the following steps. First, consider two time series $X_{t}$ and $Y_{t}$. To test if $X_{t}$ Granger-causes $Y_{t}$, we compare the following two models:
@@ 2.2 Granger Causality
2.3 Proposed Approach
In high-dimensional networks of time series, analyzing causal connectivity between specific pairs of signals can be both challenging and computationally intensive due to the complexity of modeling interactions across all nodes. The approach to be followed would be to model the whole material if there were no high dimensionality problems. However, modeling under high dimensionality requires fitting highly parameterized models to the entire network, leading to difficult-to-interpret estimates and an increased risk of overfitting. To overcome these challenges, we propose a practical and easily implementable method that focuses on isolating the nodes/channels of interest by effectively reducing the dimensionality of the network. Our approach utilizes spectral dynamic principal component analysis to capture the essential variability of the background network in a low-dimensional representation. By partialling out the influence of other nodes through sDPCA, we can isolate the COI and accurately assess the causal connectivity between them using conventional Granger causality framework. This method simplifies the analysis by avoiding the need to model the entire high-dimensional network and enhances interpretability when the primary interest is in the causal relationship between specific signals. The proposed approach is given in the Procedure outline below.
Input: High-dimensional Networks of Signals $V=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
@@ 2.3 Proposed Approach
1. Identify Nodes of Interest: NOI:$\{Y^{*}_{t},X^{*}_{t}\}$. In practice, each node can also be a cluster of nodes represented by sDPCA scores.
@@ 2. Apply sDPCA:
    Estimate spectral density matrix $F_{\mathbf{Z}}(\omega)$ for $\mathbf{Z}_{t}$.
@@
    Compute dynamic eigenvectors $\varphi_{j}(\omega)$ and eigenvalues $\lambda_{j}(\omega)$.
  3. iii
    Derive dynamic principal component filters $\phi_{k}^{(j)}$.
  4. iv
    Calculate dynamic principal component scores $\text{dpc}_{t}^{(j)}$.
3. Partial Out Influence of Other Nodes:
$Y_{t}=Y^{*}_{t}-\hat{E}\bigl{(}Y^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}$
$X_{t}=X^{*}_{t}-\hat{E}\bigl{(}X^{*}_{t}\big{|}\mathbf{p}_{t}\bigr{)}$